OpenStudy (anonymous):

Sal can paint a room in 6 hours while it takes Steve 5 hours. How long would it take them to paint the room together?

OpenStudy (saifoo.khan):

\[\frac{1}{6}+\frac{1}{5}=\frac{1}{x}\]

OpenStudy (saifoo.khan):

x = 30/11

OpenStudy (anonymous):

\[(\frac{1}{6}+\frac{1}{5})t=1\] \[\frac{11}{30}t=1\] \[t=\frac{30}{11}\]

OpenStudy (anonymous):

clear or not?

OpenStudy (anonymous):

i'm looking for a number of hours I think?

OpenStudy (anonymous):

trick is to make the job 1 for "one whole job"

OpenStudy (anonymous):

then just add the rates. one does 1/5 per hour, other does 1/6 per hour so combined they do \[\frac{1}{5}+\frac{1}{6}=\frac{11}{30}\] per hour. that is their combined rate

OpenStudy (anonymous):

you want the whole job done so set \[\frac{11}{30}t=1\] where t is time and get \[t=\frac{30}{11}\]

OpenStudy (anonymous):

okay, okay... yes that is the right answer!! I was confused. thnx!!

OpenStudy (anonymous):

if you don't like the form of the answer you could write \[2\tfrac{8}{11}\] hours

OpenStudy (anonymous):

yw